According to the IRS, a single taxpayer with a taxable income between $85,650 and $178,650 in 2012 paid $17,442, plus 28% of the amount over 85,650. If Lorinda paid 22,549 in federal income tax for 2012, then what was his taxable income in 2012
Answer by math_tutor2020(3817)


x = taxable income that is somewhere between $85,650 and $178,650
x-85650 = extra amount over $85650
28% of that is represented by 0.28*(x - 85650)

Add on 17,442 so we determine the amount of tax paid.
17442 + 0.28*(x - 85650) = tax paid
17442 + 0.28*(x - 85650) = 22549
17442 + 0.28*x - 23982 = 22549
-6540 + 0.28*x = 22549
0.28*x = 22549 + 6540
0.28*x = 29089
x = 29089/0.28
x = 103889.28571428571 approximately
x = 103889.29


Answer: $103,889.29